# Standalone 16-bit Neural Network Training: Missing Study for Hardware-Limited Deep Learning Practitioners

### v5

@@ 4.4 Results
Quantitative results are presented in Tab. 2, which outlines test accuracy and training duration with a side-by-side comparison of IEEE 16-bit Floating Point (FP16), 32-bit Floating Point (FP32), and mixed precision (MP). While it’s not obvious to single out an absolute winner, the 16-bit floating-point (FP16) clearly stands out in terms of computational speed.
One of the most striking observations is with the VGG-16 model, where FP16’s training time clocks in at a mere 377s, nearly half the time FP32 requires, 857s. This translates to a remarkable speedup of 2.2x. The efficiency of FP16 is further emphasized in Xception, with FP16 needing 324s compared to the substantial 611s that FP32 requires, again showcasing a 1.9x speedup.
When it comes to accuracy, FP16 performs competitively across most models. For example, in the AlexNet model, FP16 achieves an accuracy of 76.0%, slightly higher than FP32’s 75.8% and MP’s 75.9%. In VGG-16, FP16’s accuracy is 83.7%, which is slightly lower than FP32’s 83.9% and MP’s 83.8%. Similarly, FP16’s accuracy in ResNet-32 is 80.9%, matching FP32 and MP. In some models like DenseNet121, FP16’s accuracy is slightly lower than FP32 and MP by 0.5% and 0.9% respectively. On average, FP16’s accuracy is only 0.2% lower than FP32 and MP across the 12 examined architectures.
Despite the slight variations in accuracy, FP16 significantly improves computational training speed. On average, FP16 processes are 1.6 times faster than FP32 and 1.3 times faster than MP. This considerable increase in training speed makes FP16 an attractive option for practical applications where time efficiency is crucial.
Test accuracy comparison throughout the training process is depicted in Fig. 4. Neither FP16, FP32, nor mixed-precision emerges as a consistent frontrunner across all models. Fig. 5 offers a graphical summary of our outcomes across 50 random seeds showcased in a boxplot format. This encompasses metrics such as median, standard deviation, and outliers for the 12 architectures we evaluated. Based on median values, FP16 appears to slightly edge out the competition in VGG-16, ResNet-32, and ResNet-56. Concerning standard deviation, FP16 exhibits a smaller value across all architectures, with AlexNet being the sole exception. However, when assessing robustness, as indicated by outliers, FP16 demonstrates certain vulnerabilities, a shortcoming we recognize in Sect. 5.
In conclusion, our results indicate that utilizing IEEE 16-bit operations in CNN models and ViT significantly reduces run time without compromising accuracy, especially when compared to 32-bit and mixed precision.
5 Discussion
Although our results suggest that IEEE 16-bit floating-point can function effectively as a standalone, there are important caveats we would like to highlight.
Table 3: 16-bit (FP16) vs 32-bit (FP32) Precision Performance Comparison in Image Classification with Data Augmentation. MNIST [34] is used for DNN (Fully connected Network). CIFAR-10 Dataset [30] is used for ResNets [21] and VIT [13]. The results represent averages over experiments conducted with 4 random seeds and 300 epochs.
Model Test Accuracy Test Accuracy Acc Diff Speedup
FP16 FP32 (FP32 - FP16) (FP32 / FP16)
DNN 98.1±$\pm{0.2}$ 98.1±$\pm{0.3}$ 0.0 1.9x
ResNet-56 [21] 82.9±$\pm{0.3}$ 83.0±$\pm{0.3}$0.1 1.3x
ResNet-110 [21] 82.8±$\pm{0.3}$ 82.1±$\pm{0.4}$ 0.7 1.3x
VIT-8 [13] 74.2±$\pm{0.3}$ 74.5±$\pm{0.4}$0.3 1.4x
VIT-12 [13] 74.7±$\pm{0.3}$ 74.9±$\pm{0.3}$0.2 1.6x
5.1 Data augmentation
In our research, we initially did not include data augmentation to clearly demonstrate the effectiveness of 16-bit (FP16) precision models in straightforward image classification tasks on the MNIST dataset [34] and CIFAR-10 dataset [30]. Our main objective was to ascertain whether FP16 precision could be similar or exceed the performance of 32-bit (FP32) precision and mixed precision models. Following this, we introduced data augmentation to further test these models under more complex conditions. The experiment focuses on comparing the efficacy of 16-bit (FP16) and 32-bit (FP32) precision formats across DNN, two versions of the ResNet architecture — ResNet-56, ResNet-110 [21], and two ViT models [13].
In our experimental setup, we applied random left-right and up-down flips, adjusted the brightness by up to 10%, and rotated images within a range of -0.2 to 0.2 radians. The results, as shown in Table 3, compare FP16 and FP32 performances across different ResNet models, highlighting FP16’s consistent performance even with data augmentation. The data, averaged from experiments with four random seeds. This comprehensive approach, from non-augmented to augmented conditions, aimed to assess the viability of 16-bit precision in enhancing computational efficiency and accuracy across varied training environments.
5.2 Numerical Stability
Excluding outlier cases, the IEEE 16-bit performance is consistently on par with or better than both 32-bit and mixed precision in 4 of the 7 architectures we tested, as visualized in Figure 5. However, when considering these outliers, FP16 underperforms in terms of accuracy in 3 of the 7 architectures. This discrepancy is likely attributed to the known issues of overflow/underflow associated with the limited range of FP16. Delving into the reasons for this instability and potential countermeasures falls beyond the purview of this paper.
While such numerical instability might initially seem to deter the adoption of FP16 as a default choice, our results affirm that these anomalies are rare. Moreover, their effects can be counterbalanced by averaging results across multiple runs. This strategy can be particularly beneficial when working within strict memory constraints.
5.3 Hyperparameter Fine-tuning
As shown in Table 4, we conducted a performance analysis of the Adam optimizer [28] using float16 precision across different epsilon values. The results indicate that float16 does not perform optimally with the Adam optimizer when the epsilon parameter is set below $1\times 10^{-5}$. This suboptimal performance is particularly evident in the significant drop in test accuracy for epsilon values of $1\times 10^{-5}$ and lower.
@@ 5.3 Hyperparameter Fine-tuning
To address this issue, we explored the use of the Stochastic Gradient Descent (SGD) optimizer, which did not exhibit the same performance problems under these conditions. Consequently, hyperparameter tuning for the 16-bit neural network was unnecessary when using the SGD optimizer.
$\text{RMSProp}:w_{t}\leftarrow w_{t-1}-\eta\frac{g_{t}}{\sqrt{v_{t}}+\epsilon}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (5)
$\text{ADAM}:w_{t}\leftarrow w_{t-1}-\eta\frac{\hat{m_{t}}}{\sqrt{\hat{v_{t}}}+\epsilon}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (6)
@@
The analysis in Table 4 underscores the necessity of careful tuning of the epsilon parameter when using the Adam optimizer with float16 precision to avoid significant drops in performance.
5.4 Implementing 16-bit Batch Normalization
TensorFlow’s existing batch normalization layer doesn’t natively support 16-bit operations. Instead, it converts 16-bit inputs to 32-bit for calculations and then reverts back to 16-bit, introducing computational overhead due to the intermediate 32-bit processing. To meet our goal of evaluating standalone IEEE 16-bit neural networks, we implemented a dedicated 16-bit batch normalization layer. All our experimental results are based on this adapted layer, which fully supports 16-bit operations without requiring conversion to 32-bit.
@@ 6 Conclusion
In this study, we presented the first comprehensive theoretical and empirical analysis of standalone IEEE 16-bit floating-point neural networks. Despite the prevailing assumption that 16-bit floating-point neural networks can achieve results comparable to full (32-bit) precision, this study is the first to rigorously demonstrate this through both theoretical analysis and empirical evidence. This work fills a critical gap in the literature, providing rigorous proof that standalone IEEE 16-bit precision offers an effective and accessible option for reducing memory usage and improving computational speed, particularly for practitioners constrained by limited hardware resources.
The importance of this study is especially pronounced in resource-limited environments, where access to high-end hardware capable of supporting lower precision formats, such as FP8 or FP4, is often unavailable. Our findings show that standalone 16-bit precision can enable efficient machine learning training without the need for expensive or specialized hardware, making it a practical and valuable option for a wide range of users.
While our analysis highlights the benefits of 16-bit precision, we also identified certain limitations, particularly in relation to numerical stability in specific cases, which can affect performance. However, our theoretical formalization of floating-point errors and classification tolerance provides valuable insights into how 16-bit precision can approximate 32-bit results under the right conditions.
Overall, this work encourages broader adoption of standalone IEEE 16-bit floating-point neural networks, providing both theoretical and practical guidance. By addressing key challenges and demonstrating the feasibility of 16-bit precision in resource-limited settings, we believe this research opens up new possibilities for more accessible and efficient machine learning across diverse environments.
@@ 7 Statements and Declarations
The researcher claims no conflicts of interest. The datasets generated and/or analyzed during the current study, as well as the code for reproducing the results, are available on GitHub at https://github.com/YUNBLAK/standalone_16bits_nn.
Funding
This research was partially supported by the National Research Foundation of Korea (NRF) under Grant RS-2022-00165647. This work was also supported by the Innovative Human Resource Development for Local Intellectualization program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-RS-2023-00259678).

### v6

@@ 4.3 Results
As summarized in Tab. 2, our results demonstrate that IEEE 16-bit floating point (FP16) offers a significant computational advantage over both 32-bit (FP32) and mixed precision (MP), with an average speedup of 1.6$\times$ over FP32 and 1.3$\times$ over MP. For instance, FP16 reduces VGG-16 training time from 857s (FP32) to 377s, and Xception from 611s to 324s. In terms of accuracy, FP16 performs competitively, with only a 0.2% average decrease compared to FP32 and MP across 12 architectures. In some cases, such as AlexNet and ResNet-32, FP16 matches or even slightly outperforms the alternatives. Fig. 3 illustrates test accuracy trends over training epochs, showing no consistent winner across precisions. The boxplot in Fig. 4, averaged over 50 random seeds, further shows that FP16 yields similar or better median accuracy with lower standard deviation in most models, though occasional outliers suggest reduced robustness, as discussed in Sect. 5. Overall, FP16 proves to be an efficient and accurate alternative for deep learning under hardware constraints.
5 Discussion: Numerical Stability and Hyperparameter Tuning
While FP16 performs competitively across most architectures, we observe occasional performance drops in a few models due to numerical instability, particularly from overflow or underflow errors inherent to the limited dynamic range of 16-bit floats. As shown in Fig. 4, FP16 underperforms in 3 out of 7 architectures when considering outlier cases. However, these anomalies are rare and can be mitigated by averaging results across multiple runs, making FP16 a viable option even under strict memory constraints.
@@ 5 Discussion: Numerical Stability and Hyperparameter Tuning
We further analyze the impact of optimizer hyperparameters in FP16 training. Table 3 presents a performance comparison using the Adam optimizer [24] on CIFAR-10 [26] with varying epsilon values. The results show that FP16 suffers from significant accuracy degradation when $\epsilon<10^{-5}$, likely due to instability in the denominator of the update rule:
@@
A similar concern applies to RMSProp [20]. The default $\epsilon=10^{-7}$ in TensorFlow, while suitable for FP32, may cause numerical issues in FP16 when $\hat{v_{t}}$ or $v_{t}$ are close to zero. To address this, we adopted SGD [14], which exhibited stable performance without requiring tuning. This suggests that FP16 networks can avoid instability by using simpler optimizers and avoiding extreme hyperparameter values.
@@ 6 Conclusion
This study presents the first comprehensive theoretical and empirical analysis of standalone IEEE 16-bit floating-point neural networks. Contrary to the common reliance on mixed or higher precision, we rigorously demonstrate that 16-bit precision alone can match 32-bit performance in both accuracy and stability across a range of models, while offering substantial improvements in training speed and memory efficiency. These findings are particularly relevant for practitioners with limited hardware access, where support for formats like FP8 or FP4 is unavailable. Our theoretical framework based on floating-point error and classification tolerance provides insights into when 16-bit can safely replace 32-bit precision. While some numerical instability remains in edge cases, our results support broader adoption of 16-bit training as a practical, efficient, and accessible solution for deep learning in resource-constrained environments.
@@ 7 Statements and Declarations
The researcher claims no conflicts of interest. This research was partially supported by the National Research Foundation of Korea (NRF) under Grant RS-2022-00165647. This work was also supported by the Innovative Human Resource Development for Local Intellectualization program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-RS-2023-00259678).